OpenStudy (anonymous):

Which equation would best help solve the following problem? Adina has a rectangular garden that measures 9 m wide by 13 m long. She wants to increase the area to 192 m2 by increasing the width and length by the same amount. What will be the dimensions of the new garden? A. 192 = (9)(13) B. 192 = (9)(13 + x2) C. 192 = (9 – x)(13 – x) D. 192 = (9 + x)(13 + x)

OpenStudy (anonymous):

@animal108

OpenStudy (anonymous):

Area = length*breadth Let the increment in the sides be x then new length = 13+x new breadth = 9 +x new area = 192 so 192 = (13+x)*(9+x)

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

nvm i got it